# Beyond One-Size-Fits-All: Tailored Benchmarks for Efficient Evaluation

### v1

@@ Abstract
Evaluating models on large benchmarks is very resource-intensive, especially during the period of rapid model evolution. Existing efficient evaluation methods estimate the performance of target models by testing them only on a small and static coreset of the benchmark, which is derived from the publicly available evaluation results of source models. These methods rely on the assumption that target models have high prediction consistency with source models. However, we demonstrate that it doesn’t generalize well in practice. To alleviate the inconsistency issue, we present TailoredBench, a method that conducts customized evaluation tailored to each target model. Specifically, a Global-coreset is first constructed as a probe to identify the most consistent source models for each target model with an adaptive source model selection strategy. Afterwards, a scalable K-Medoids clustering algorithm is proposed to extend the Global-coreset to a tailored Native-coreset for each target model. According to the predictions on Native-coresets, we obtain the performance of target models on the whole benchmark with a calibrated estimation strategy. Comprehensive experiments on 5 benchmarks across over 300 models demonstrate that compared to best performing baselines, TailoredBench achieves an average reduction of 31.4% in MAE of accuracy estimates under the same inference budgets, showcasing strong effectiveness and generalizability.
@@ 1 Introduction
To achieve efficient evaluation, some studies (Vivek et al., 2024; Polo et al., 2024) have explored the following paradigm: step 1. constructing example embeddings according to the predictions from a set of source models (which are freely available for popular leaderboards***https://huggingface.co/open-llm-leaderboard,https://rank.opencompass.org.cn,https://crfm.stanford.edu/helm); step 2. clustering the benchmark and selecting the cluster centroids to form a coreset (typically less than 100 examples); step 3. approximating the performance of target models under evaluation based on their predictions on the coreset. (See detailed related works in Appendix A.) Underlying this approach is the assumption that performance patterns generalize: if source models respond similarly to two examples $a$ and $b$, then a target model’s performance on $a$ can be used to estimate its performance on $b$.
@@ 3.2 Main Results
Tables 1 present a comprehensive comparison between our TailoredBench method and the best baseline approaches for each metric across all benchmarks. Full results are available in Appendix B.1. In our experiments, we allocated 10 examples to the G-set and averaged the outcomes over 100 randomized trials to ensure statistical reliability. The inference count—defined as the number of examples in the N-set for our method—varied from 20 to 40.
As demonstrated in the tables, our method consistently outperforms baseline approaches in both Kendall’s $\tau$ and MAE metrics across all inference counts and benchmarks featuring different correctness types. When the inference count increases, the performance of our method continues to improve, evidenced by a steady increase in Kendall’s $\tau$ and a continuous decrease in MAE. Notably, compared to best performing baselines, our approach achieves nearly a 31.4% reduction in MAE. These results indicate that our method effectively estimates the relative performance among target models and provides more accurate estimations of their performance on the entire benchmark. Furthermore, compared to the static AnchorPoints method, our approach significantly improves both Kendall’s $\tau$ and MAE metrics, highlighting its effectiveness in adaptively selecting a more representative N-set for each target model and thereby improving estimation accuracy. We also calculate the accuracy of our method in ranking the performance between every pair of target models. The results show that the accuracy reached 96.0% on the Hellaswag benchmark and 93.6% on the GSM8K benchmark. In terms of robustness, Appendix B.8 demonstrates that our method exhibits significantly lower variance compared to the baselines.
@@ Appendix A Related Works
Coreset Selection for Efficient Benchmarking:
As LLMs proliferate and version updates accelerate, the cost of thoroughly evaluating each model across all benchmarks has become prohibitive, leading to methods that subsample the most representative subsets from each benchmark for more efficient evaluation. Vivek et al. (2024) clusters examples directly using the confidence scores provided by source models, leveraging these scores to select an optimal subset. Similarly, Polo et al. (2024) employs an Item Response Theory (IRT) model, trained on the success matrix of each source model across various examples, to derive the latent representations of examples for clustering. Pacchiardi et al. (2024) introduces a generic assessor framework that predicts the performance of a new LLM on unseen instances using its results on a small reference set, achieving comparable accuracy to full-scale evaluations. Perlitz et al. (2023) proposes Flash-HELM, which dynamically adjusts the sizes of randomly selected subsets based on model ranking, where higher-ranked models are evaluated with greater precision. Prabhu et al. (2024) proposes the Sort & Search (S&S) strategy, which leverages the difficulties of examples and dynamic programming to select the coreset. Xu et al. (2024) synthesizes several methods and dynamically chooses the optimal subset selection method for each benchmark but requires many examples to determine the best approach. Despite these advancements, these methods often struggle with substantial distribution shifts between the source and target models, caused by the discrepancy between their predictive consistency, potentially causing significant distortion in estimating the target model’s performance. Extending the approach of Vivek et al. (2024), our work alleviates this issue by dynamically selecting a native source model set with the highest prediction consistency to the target model, ensuring the selection of a tailored coreset for each target model that best represents the benchmark.

### v2

@@ Abstract
Evaluating models on large benchmarks is very resource-intensive, especially during the period of rapid model evolution. Existing efficient evaluation methods estimate the performance of target models by testing them only on a small and static coreset of the benchmark, which is derived from the publicly available evaluation results of source models. These methods rely on the assumption that target models have high prediction consistency with source models. However, we demonstrate that it doesn’t generalize well in practice. To alleviate the inconsistency issue, we present TailoredBench, a method that conducts customized evaluation tailored to each target model. Specifically, a Global-coreset is first constructed as a probe to identify the most consistent source models for each target model with an adaptive source model selection strategy. Afterwards, a scalable K-Medoids clustering algorithm is proposed to extend the Global-coreset to a tailored Native-coreset for each target model. According to the predictions on Native-coresets, we obtain the performance of target models on the whole benchmark with a calibrated estimation strategy. Comprehensive experiments on 5 benchmarks across over 300 models demonstrate that compared to best performing baselines, TailoredBench achieves an average reduction of 31.4% in MAE of accuracy estimates under the same inference budgets, showcasing strong effectiveness and generalizability [^1] Our code is available at https://github.com/marvelcell/TailoredBench.
@@ 1 Introduction
To achieve efficient evaluation, some studies (Vivek et al., 2024; Polo et al., 2024) have explored the following paradigm: step 1. constructing example embeddings according to the predictions from a set of source models (which are freely available for popular leaderboardshttps://huggingface.co/open-llm-leaderboard,https://rank.opencompass.org.cn,§§§https://crfm.stanford.edu/helm); step 2. clustering the benchmark and selecting the cluster centroids to form a coreset (typically less than 100 examples); step 3. approximating the performance of target models under evaluation based on their predictions on the coreset. Underlying this approach is the assumption that performance patterns generalize: if source models respond similarly to two examples $a$ and $b$, then a target model’s performance on $a$ can be used to estimate its performance on $b$.
@@
2 Related Works
As LLMs proliferate and version updates accelerate, the cost of thoroughly evaluating each model across all benchmarks has become prohibitive, leading to methods that subsample the most representative subsets from each benchmark for more efficient evaluation. Vivek et al. (2024) clusters examples directly using the confidence scores provided by source models, leveraging these scores to select an optimal subset. Similarly, Polo et al. (2024) employs an Item Response Theory (IRT) model, trained on the success matrix of each source model across various examples, to derive the latent representations of examples for clustering. Pacchiardi et al. (2024) introduces a generic assessor framework that predicts the performance of a new LLM on unseen instances using its results on a small reference set, achieving comparable accuracy to full-scale evaluations. Perlitz et al. (2023) proposes Flash-HELM, which dynamically adjusts the sizes of randomly selected subsets based on model ranking, where higher-ranked models are evaluated with greater precision. Prabhu et al. (2024) proposes the Sort & Search (S&S) strategy, which leverages the difficulties of examples and dynamic programming to select the coreset. (See more related works in Appendix A.)
@@ 4.2 Main Results
Table 1 present a comprehensive comparison between our TailoredBench method and the best baseline approaches for each metric across all benchmarks. Full results are available in Appendix B.1. In our experiments, we allocated 10 examples to the G-set and averaged the outcomes over 100 randomized trials to ensure statistical reliability. The inference count—defined as the number of examples in the N-set for our method—varied from 20 to 40.
As demonstrated in the table, our method consistently outperforms baseline approaches in both Kendall’s $\tau$ and MAE metrics across all inference counts and benchmarks featuring different correctness types. When the inference count increases, the performance of our method continues to improve, evidenced by a steady increase in Kendall’s $\tau$ and a continuous decrease in MAE. Notably, compared to best performing baselines, our approach achieves nearly a 31.4% reduction in MAE. These results indicate that our method effectively estimates the relative performance among target models and provides more accurate estimations of their performance on the entire benchmark. Furthermore, compared to the static AnchorPoints method, our approach significantly improves both Kendall’s $\tau$ and MAE metrics, highlighting its effectiveness in adaptively selecting a more representative N-set for each target model and thereby improving estimation accuracy. We also calculate the accuracy of our method in ranking the performance between every pair of target models. The results show that the accuracy reached 96.0% on the Hellaswag benchmark and 93.6% on the GSM8K benchmark. In terms of robustness, Appendix B.8 demonstrates that our method exhibits significantly lower variance compared to the baselines.
@@ 5 Conclusions
Acknowledgments
This work is supported by Beijing Natural Science Foundation (No.4222037, L181010).
@@ Appendix A Related Works
Xu et al. (2024) synthesizes several methods and dynamically chooses the optimal subset selection method for each benchmark but requires many examples to determine the best approach. Despite these advancements, these methods often struggle with substantial distribution shifts between the source and target models, caused by the discrepancy between their predictive consistency, potentially causing significant distortion in estimating the target model’s performance. Extending the approach of Vivek et al. (2024), our work alleviates this issue by dynamically selecting a native source model set with the highest prediction consistency to the target model, ensuring the selection of a tailored coreset for each target model that best represents the benchmark.